OpenStudy (anonymous):

Can someone tell me how to graph How do I graph 3x-2y = -6 using slope intercept method? I'm having trouble getting it in that method

OpenStudy (anonymous):

y = mx + b

OpenStudy (anonymous):

0, b plot that m = slope = rise/run connect those two points....

OpenStudy (anonymous):

For some reason I can't get it into y=mx+b to plot it. I'm being told to subtract something from both sides, which I did but I came up with a decimal...

OpenStudy (anonymous):

you are just using algebra to isolate the y

OpenStudy (anonymous):

 do not put your slope into decimal...leave it this way rise over run effect... no go back to my steps earlier do plot accordingly